OpenStudy (anonymous):

fraction arithmetic 1. 6/7 x 5/3 2. 4 2/5 x (-3 1/3) 3. 2 2/5 divide (-1 1/2) 4. - 8/9 divide (- 4/3) 5. 5/8 x (- 4/15) word problem 6. My recipe calls for 2/3 cups of white flour and 2 1/2 cups of whole wheat flour. How much flour do i need in total for my recipe?

OpenStudy (anonymous):

Please give a good answer so i get it step by step each problem thanks

OpenStudy (curry):

3 1/6 = problem #6

OpenStudy (curry):

add 2/3 to 2 1/2 = 4/6 + 21/6

OpenStudy (anonymous):

step by step

OpenStudy (anonymous):

thanks thats for # 6 right

OpenStudy (curry):

1 3/7 = problem #1 please be my fan 5*6= 30 7*3= 21

OpenStudy (anonymous):

?